### v2

@@ 4.4 Interaction in terms of pseudopotentials
A four-fermion interaction term is formed by contracting these paring operators with its conjugate
@@
[…]
@@
$C^{l}_{m_{1}m_{2}m_{3}m_{4}}=\delta_{m_{1}+m_{2},m_{3}+m_{4}}\langle sm_{1},sm_{2}|(2s-l)(m_{1}+m_{2})\rangle\langle sm_{3},sm_{4}|(2s-l)(m_{3}+m_{4})\rangle.$ (30)
The coupling strengths $U_{l}$ of the spin-$(2s-l)$ channel are called the Haldane pseudopotentials.
We also need to consider the constraint that the two fermions must be antisymmetrised : for even $l$, the orbital index is symmetrised, so the spin index must be antisymmetrised, so the two fermions form a spin-singlet which is invariant under the $\mathrm{SU}(2)$ transformation ; for odd $l$, the orbital index is antisymmetrised, so the spin index is symmetrised, breaking the flavour symmetry from $\mathrm{SU}(2)$ to $\mathbb{Z}_{2}$. Hence, an odd-$l$ pseudopotential must be added (This fact escapes the construction by density-density interaction).
@@
Each pseudopotential corresponds to a profile of interaction potential functions. The conversion between the pseudopotentials $U_{l}$ and the Legendre expansion coefficients of the potential function $\tilde{U}_{l}$
@@
where {$\ \bullet&\bullet&\bullet\end{smallmatrix}\right\}$ is the $6j$-symbol. Specifically, a local interaction $\delta(\mathbf{r}_{12})$ contains only pseudopotential $U_{0}$ ; a superlocal interaction of form $(\nabla^{2})^{l}\delta(\mathbf{r}_{12})$ contains $U_{0},U_{1},\dots,U_{l}$. Here we explicitly give the expressions for the lowest pseudopotentials
@@
More details are given in Ref. [15]. We also note that the construction for interaction Hamiltonian for LLL with pseudopotentials is not restricted to sphere. E.g., on a torus, the interaction can also be parametrise by pseudopotentials [98, 99, 100]. More details are presented in Appendix B.
For systems with more complicated continuous flavour symmetries, classification in terms of representation of flavour symmetry must also be considered, and the indices must be overall antisymmetrised. We explain that through the example of a $2N$-flavour system with $\mathrm{Sp}(N)$ symmetry [16]. The maximal flavour symmetry is $\mathrm{SU}(2N)$, so interactions must be added to break the symmetry from $\mathrm{SU}(2N)$ to $\mathrm{Sp}(N)$. The fermion operators $c_{ma}$ live in the $\mathrm{Sp}(N)$ fundamental representation, where $a$ is the $\mathrm{Sp}(N)$ index. We shall show that all the allowed terms are
@@
To find out all the four-fermion interactions allowed by the rotation symmetry $\mathrm{SO}(3)$ and flavour symmetry $\mathrm{Sp}(N)$, we classify all the fermion bilinears $c_{m_{1}a}c_{m_{2}b}$ into irreps of $\mathrm{SO}(3)\times\mathrm{Sp}(N)$. For each irrep, by contracting the bilinear with its Hermitian conjugate, we obtain an allowed four-fermion interaction term. Each fermion carries $\mathrm{SO}(3)$ spin-$s$ and $\mathrm{Sp}(N)$ fundamental. For the rotation symmetry $\mathrm{SO}(3)$, the bilinears can carry spin-$(2s-l)(l=0,\dots,2s)$ represetation ; for even $l$, the orbital indices are symmetrised ; for odd $l$, the orbital indices are antisymmetrised. For the flavour symmetry $\mathrm{Sp}(N)$, the bilinears can carry singlet $S$, traceless antisymmetric rank-2 tensor $A$ and symmetric rank-2 tensor $T$ representation ; for $S$ and $A$, the flavour indices are antisymmetrised ; for $T$, the flavour indices are symmetrised. As the two fermions altogether should be antisymmetrised, the allowed combinations are
@@
    $\mathrm{Sp}(N)$ singlet and $\mathrm{SO}(3)$ spin-$(2s-l)$ with even $l$, the bilinears are
@@
    The corresponding interaction term $H_{S,l}=\sum_{m}\Delta_{lm}^{\dagger}\Delta_{lm}$ is the even-$l$ pseudopotential for the $V$-term.
@@
    $\mathrm{Sp}(N)$ antisymmetric and $\mathrm{SO}(3)$ spin-$(2s-l)$ with even $l$, the bilinears are
@@
    The corresponding interaction term $H_{A,l}=\sum_{m}\Delta_{lm,[ab]}^{\dagger}\Delta_{lm,[ab]}$ is the even-$l$ pseudopotential for the $U$-term.
@@
    $\mathrm{Sp}(N)$ symmetric and $\mathrm{SO}(3)$ spin-$(2s-l)$ with odd $l$, the bilinears are
@@
    The corresponding interaction term $H_{T,l}=\sum_{m}\Delta_{lm,(ab)}^{\dagger}\Delta_{lm,(ab)}$ is the odd-$l$ pseudopotential for the $U$-term.
In summary, all allowed interactions are the $U_{l}$ terms with both even and odd $l$, and the $V_{l}$ terms with only even $l$.
@@ 4.5 Operator spectrum and search for conformal point
where $E_{0}$ is the ground state energy, $R$ is the radius of the sphere (here we take $R=\sqrt{N_{m}}$), and $v$ is a model-dependent speed of light. The constant $v/R$ can be determined through a calibration process, i.e., comparing the spectrum to some known properties of a CFT spectrum. The criteria to determine the conformal symmetry include
@@
    The existence of a conserved flavour symmetry current $J^{\mu}$ if there is a continuous flavour symmetry. The symmetry current typically lives in the antisymmetric rank-2 tensor representation of the flavour symmetry. E.g., if the flavour symmetry is $\mathrm{U}(1)$, then the symmetry current has charge-0 ; if the flavour symmetry is $\mathrm{O}(3)$, then the symmetry current has spin-1 and is odd under the improper $\mathbb{Z}_{2}$ transformation ; if the flavour symmetry is $\mathrm{O}(n)$ ($n\geq 4$) or $\mathrm{SU}(n)$ ($n\geq 3$), then the symmetry current lives in the antisymmetric rank-$2$ tensor representation.
@@
    where $\Box=\partial_{\mu}\partial^{\nu}$. For a spinning primary $\Phi^{\mu_{1}\dots\mu_{s}}$, its descendants has the two forms :
    $\Box^{n}\partial^{\nu_{1}}\dots\partial^{\nu_{m}}\partial_{\rho_{1}}\dots\partial_{\rho_{k}}\Phi^{\rho_{1}\dots\rho_{k}\mu_{1}\dots\mu_{s-k}}\qquad(k=0,\dots,s,\ n,m=0,1,\dots)$
@@
    $\Box^{n}\partial^{\nu_{1}}\dots\partial^{\nu_{m}}\partial_{\rho_{1}}\dots\partial_{\rho_{k}}\epsilon^{\sigma}{}_{\tilde{\mu}\tilde{\nu}}\partial^{\tilde{\nu}}\Phi^{\rho_{1}\dots\rho_{k}\tilde{\mu}\mu_{1}\dots\mu_{s-k-1}}\qquad(k=0,\dots,s-1,\ n,m=0,1,\dots)$
@@
    For the second form, the fully antisymmetric tensor $\epsilon$ alters the parity.
@@
Alternatively, one can define a cost function that depends on the tuning parameter and the speed of light and compares the scaling dimensions obtained from the fuzzy sphere and the prediction by conformal symmetry. E.g., for the Ising CFT, the tuning parameters are the pseudopotentials $\{U_{i}\}$ and the transverse field $h$. The criteria for conformal symmetry we use include the stress tensor $T^{\mu\nu}$ and the descendants $\partial^{\mu}\sigma$, $\partial^{\mu}\partial^{\nu}\sigma$, $\Box\sigma$, $\partial^{\mu}\epsilon$. The cost function is the root-mean-square of the deviations of these criteria from the expectation of the conformal symmetry
$Q^{2}(\{U_{i}\},h,v;N_{m})=\frac{1}{N_{s}}\left[(\Delta_{T^{\mu\nu}}^{(\textrm{FS})}-3)^{2}+(\Delta_{\partial^{\mu}\sigma}^{(\textrm{FS})}-\Delta_{\sigma}^{(\textrm{FS})}-1)^{2}\right.\\ \left.+(\Delta_{\partial^{\mu}\partial^{\nu}\sigma}^{(\textrm{FS})}-\Delta_{\sigma}^{(\textrm{FS})}-1)^{2}+(\Delta_{\Box\sigma}^{(\textrm{FS})}-\Delta_{\sigma}^{(\textrm{FS})}-1)^{2}+(\Delta_{\partial^{\mu}\epsilon}^{(\textrm{FS})}-\Delta_{\epsilon}^{(\textrm{FS})}-1)^{2}\right]$ (40)
@@
$\Delta_{\Phi}^{(\textrm{FS})}(\{U_{i}\},h,v;N_{m})=\frac{E_{\Phi}-E_{0}}{v/R}.$ (41)
@@ 4.6 Local observables
We have introduced how to determine the scaling dimensions from the energy spectrum. Beyond that, evaluating other CFT quantities requires realising local CFT operators on the fuzzy sphere. Any gapless local observables $\mathscr{O}(\hat{\mathbf{n}})$ on the fuzzy sphere can be written as the linear combination of CFT operators that live in the same representation of flavour symmetry and parity [^8] The realisation of CFT operators in the microscopic model has been investigated in more detail in 2d in Ref. [101]. [2]
$\mathscr{O}(\hat{\mathbf{n}},\tau)=\sum_{\alpha}\lambda_{\alpha}\Phi^{(\textrm{cyc.})}_{\alpha}(\hat{\mathbf{n}},\tau).$ (42)
Here special care should be taken for the CFT operator $\Phi^{(\textrm{cyc.})}_{\alpha}(\hat{\mathbf{n}},\tau)$ on the cylinder. A conformal transformation produces a scale factor $\Lambda(\mathbf{r})^{\Delta_{\Phi}}$ to a primary operator $\Phi$. The scale factor is $\Lambda(\mathbf{r})=r/R$ for the Weyl transformation from the flat spacetime to the cylinder. Hence [^9] We need to clarify some of the notations : $\mathscr{O}$ represents an operator in the microscopic model, and $\Phi$ represents a CFT operator ; the arguments $\Phi(\hat{\mathbf{n}})$ or $\Phi(\hat{\mathbf{n}},\tau)$ by default mean the operator is defined on a cylinder, and $\Phi(\mathbf{r})$ by default on flat spacetime.,
$\Phi^{(\textrm{cyc.})}_{\alpha}(\hat{\mathbf{n}},\tau)=\left(\frac{e^{\tau/R}}{R}\right)^{\Delta_{\Phi_{\alpha}}}\Phi_{\alpha}^{\textrm{(flat)}}(\mathbf{r}).$ (43)
For descendants, certain other factors may be produced, but the conversion factors still scale with the radius of the sphere as $R^{-\Delta}$ where $\Delta$ is the scaling dimension of the descendants. For simplicity, hereafter we focus on the equal-time correlators with $\tau=0$, for which $\Phi^{(\textrm{cyc.})}_{\alpha}(\hat{\mathbf{n}})=R^{-\Delta_{\Phi_{\alpha}}}\Phi_{\alpha}^{\textrm{(flat)}}(\mathbf{r})$. The operator with larger system size decays faster when increasing system size.
The simplest local observable is the density operator defined in Eq. (15), and its decomposition into angular modes is given in Eqs. (17) and (18). From the CFT perspective, the density operators are the superpositions of scaling operators with corresponding quantum numbers, i.e., with the same representation under flavour symmetry and parity.
@@
$n^{x}(\hat{\mathbf{n}})$ $=\lambda_{0}+\lambda_{\epsilon}\epsilon(\hat{\mathbf{n}})+\lambda_{\partial^{\mu}\epsilon}\partial^{\mu}\epsilon(\hat{\mathbf{n}})+\lambda_{T^{\mu\nu}}T^{\mu\nu}(\hat{\mathbf{n}})+\dots$ $\epsilon_{\textrm{FS}}$ $=\frac{n^{x}-\lambda_{0}}{\lambda_{\epsilon}}+\dots$
$n^{z}(\hat{\mathbf{n}})$ $=\lambda_{\sigma}\sigma(\hat{\mathbf{n}})+\lambda_{\partial^{\mu}\sigma}\partial^{\mu}\epsilon(\hat{\mathbf{n}})+\lambda_{\partial^{\mu}\partial^{\nu}\sigma}\partial^{\mu}\partial^{\nu}\sigma(\hat{\mathbf{n}})+\dots$ $\sigma_{\textrm{FS}}$ $=\frac{n^{z}}{\lambda_{\sigma}}+\dots$ (44)
@@
We first consider the insertion of a single operator $\langle\Phi_{1}|\Phi_{2}(\hat{\mathbf{n}})|\Phi_{3}\rangle$. It helps us produce the OPE coefficients. For the simplest example of three scalars,
$f_{\Phi_{1}\Phi_{2}\Phi_{3}}=\lim_{r_{\infty}\to\infty}r_{\infty}^{-2\Delta_{\Phi_{1}}}\langle\Phi_{1}(x_{\infty})\Phi_{2}(x)\Phi_{3}(0)\rangle_{\textrm{flat}}=\langle\Phi_{1}|\Phi_{2}^{(\textrm{flat})}(x)|\Phi_{3}\rangle$ (45)
@@
After the Weyl transformation from the flat spacetime to the cylinder, we obtain the expression on the fuzzy sphere
$f_{\Phi_{1}\Phi_{2}\Phi_{3}}=R^{\Delta_{\Phi_{2}}}\langle\Phi_{1}|\Phi^{(\textrm{cyl.})}_{2}(\hat{\mathbf{n}})|\Phi_{3}\rangle.$ (48)
The UV realisation of $\Phi_{2}$ contains many other operators with different spins. By integrating the correlation function against different spherical harmonics, i.e., take the angular modes of the operator inserted
$\int\mathrm{d}^{2}\hat{\mathbf{n}}\,\bar{Y}_{lm}(\hat{\mathbf{n}})\langle\Phi_{1}|\Phi_{2}(\hat{\mathbf{n}})|\Phi_{3}\rangle=\langle\Phi_{1}|\Phi_{2,lm}|\Phi_{3}\rangle,$ (49)
we can filter out the subleading contributions with different spins. For the spinning operators, this also tells us about different OPE structures. By taking $\Phi_{3}=\mathbb{I}$, we can recover the 2-pt functions
@@
$\Phi_{2}(\hat{\mathbf{n}})|0\rangle$ $=R^{-\Phi_{2}}\left[|\Phi_{2}\rangle+\lambda^{\prime}_{\mu}(\hat{\mathbf{n}})|\partial^{\mu}\Phi_{2}\rangle+\lambda^{\prime\prime}(\hat{\mathbf{n}})|\Box\Phi_{2}\rangle+\lambda^{\prime\prime}_{\mu\nu}(\hat{\mathbf{n}})|\partial^{\mu}\partial^{\nu}\Phi_{2}\rangle\right].$ (50)
It is worth noting that acting a primary $\Phi_{2}(\hat{\mathbf{n}})$ on the vacuum also produces various descendants in the multiplet.
@@
$\lambda_{0}=\frac{R^{-2}}{\sqrt{4\pi}}\langle 0|n^{x}_{00}|0\rangle,\quad\lambda_{\epsilon}=\frac{R^{\Delta_{\epsilon}-2}}{\sqrt{4\pi}}\langle\epsilon|n^{x}_{00}|0\rangle,\quad\lambda_{\sigma}=\frac{R^{\Delta_{\sigma}-2}}{\sqrt{4\pi}}\langle\sigma|n^{z}_{00}|0\rangle.$ (51)
@@
$f_{\sigma\sigma\epsilon}$ $=R^{\Delta_{\sigma}}\langle\epsilon|\sigma(\hat{\mathbf{n}})|\sigma\rangle=\frac{\langle\epsilon|n_{00}^{z}|\sigma\rangle}{\langle 0|n_{00}^{z}|\sigma\rangle}+\mathscr{O}(R^{-2})$
$=R^{\Delta_{\epsilon}}\langle\sigma|\epsilon(\hat{\mathbf{n}})|\sigma\rangle=\frac{\langle\sigma|n^{x}_{00}|\sigma\rangle-\langle 0|n^{x}_{00}|0\rangle}{\langle\epsilon|n^{x}_{00}|0\rangle}+\mathscr{O}(R^{-(3-\Delta_{\epsilon})}).$ (52)
For the first line, the subleading contribution comes from the contribution of the descendant $\Box\sigma$ to $n_{00}^{z}$. As $\sigma(\hat{\mathbf{n}})$ scales as $R^{-\Delta_{\sigma}}$ and $\Box\sigma(\hat{\mathbf{n}})$ as $R^{-\Delta_{\sigma}-2}$,
@@
Here $c_{1}$ and $c^{\prime}_{1}$ are constant factors that represent the contribution of $\Box\sigma$ and do not scale with system size. Hence, the subleading contribution scales as $R^{-2}$. For the second line, the subleading contribution comes from the stress tensor $T^{\mu\nu}$. Similarly, the power of the scaling is the difference of the scaling dimension $R^{-(\Delta_{T^{\mu\nu}}-\Delta_{\epsilon})}=R^{-(3-\Delta_{\epsilon})}$.

### v3

@@ 4.4 Interaction in Terms of Pseudo-Potentials
A four-fermion interaction term is formed by contracting these pairing operators with its conjugate
@@
$H=\sum_{l,m_{1}m_{2}m_{3}m_{4}}U_{l}C^{l}_{m_{1}m_{2}m_{3}m_{4}}c^{\dagger}_{m_{1}\uparrow}c^{\dagger}_{m_{2}\downarrow}c_{m_{3}\downarrow}c_{m_{4}\uparrow}-h\sum_{m}(c_{m\uparrow}^{\dagger}c_{m\downarrow}+\text{h.\penalty 10000\ c.}),$ (28)
@@
$C^{l}_{m_{1}m_{2}m_{3}m_{4}}=\delta_{m_{1}+m_{2},m_{3}+m_{4}}\langle sm_{1},sm_{2}|(2s-l)(m_{1}+m_{2})\rangle\langle(2s-l)(m_{3}+m_{4})|sm_{3},sm_{4}\rangle.$ (29)
The coupling strengths $U_{l}$ of the spin-(2sl)(2s-l) channel are called the Haldane pseudo-potentials.
We also need to consider the constraint that the two fermions must be anti-symmetrised : for even $l$, the orbital index is symmetrised, so the spin index must be anti-symmetrised, so the two fermions form a spin-singlet which is invariant under the $\mathrm{SU}(2)$ transformation ; for odd $l$, the orbital index is anti-symmetrised, so the spin index is symmetrised, breaking the flavour symmetry from $\mathrm{SU}(2)$ to $\mathbb{Z}_{2}$. Hence, an odd-$l$ pseudo-potential must be added (This fact escapes the construction by density-density interaction).
@@
Each pseudo-potential corresponds to a profile of interaction potential functions. The conversion between the pseudo-potentials $U_{l}$ and the Legendre expansion coefficients of the potential function $\tilde{U}_{l}$
@@
where $\left\{\begin{smallmatrix}\bullet$$\bullet&\bullet\\$ $\bullet$$\bullet&\bullet\end{smallmatrix}\right\}$ is the $6j$-symbol. Specifically, a local interaction $\delta(\mathbf{r}_{12})$ contains only pseudo-potential $U_{0}$ ; a superlocal interaction of form $(\nabla^{2})^{l}\delta(\mathbf{r}_{12})$ contains $U_{0},U_{1},\dots,U_{l}$. Here we explicitly give the expressions for the lowest pseudo-potentials
@@
More details are given in Ref. [15]. We also note that the construction for interaction Hamiltonian for LLL with pseudo-potentials is not restricted to sphere. E. g., on a torus, the interaction can also be parametrise by pseudo-potentials [142, 143, 144]. More details are presented in Appendix B.
For systems with more complicated continuous flavour symmetries, classification in terms of representation of flavour symmetry must also be considered, and the indices must be overall anti-symmetrised. We explain that through the example of a $2N$-flavour system with $\mathrm{Sp}(N)$ symmetry [16]. The maximal flavour symmetry is $\mathrm{SU}(2N)$, so interactions must be added to break the symmetry from $\mathrm{SU}(2N)$ to $\mathrm{Sp}(N)$. The fermion operators $c_{ma}$ live in the $\mathrm{Sp}(N)$ fundamental representation, where $a$ is the $\mathrm{Sp}(N)$ index. We shall show that all the allowed terms are
@@
[…]
@@
    $\mathrm{Sp}(N)$ singlet and $\mathrm{SO}(3)$ spin-(2sl)(2s-l) with even $l$, the bilinears are
@@
    The corresponding interaction term $H_{S,l}=\sum_{m}\Delta_{lm}^{\dagger}\Delta_{lm}$ is the even-$l$ pseudo-potential for the $V$-term.
@@
    $\mathrm{Sp}(N)$ anti-symmetric and $\mathrm{SO}(3)$ spin-(2sl)(2s-l) with even $l$, the bilinears are
@@
    The corresponding interaction term $H_{A,l}=\sum_{m}\Delta_{lm,[ab]}^{\dagger}\Delta_{lm,[ab]}$ is the even-$l$ pseudo-potential for the $U$-term.
@@
    $\mathrm{Sp}(N)$ symmetric and $\mathrm{SO}(3)$ spin-(2sl)(2s-l) with odd $l$, the bilinears are
@@
    The corresponding interaction term $H_{T,l}=\sum_{m}\Delta_{lm,(ab)}^{\dagger}\Delta_{lm,(ab)}$ is the odd-$l$ pseudo-potential for the $U$-term.
In summary, by classifying all the bilinears, we show that all allowed interactions are the $U_{l}$ terms with both even and odd $l$, and the $V_{l}$ terms with only even $l$.
@@ 4.5 Operator Spectrum and Search for Conformal Point
where $E_{0}$ is the ground state energy, $R$ is the radius of the sphere (here we take $R=\sqrt{N_{m}}$), and $v$ is a model-dependent speed of light. The constant $v/R$ can be determined through a calibration process, i. e. comparing the spectrum to some known properties of a CFT spectrum. The criteria to determine the conformal symmetry include
@@
    The existence of a conserved flavour symmetry current $J^{\mu}$ if there is a continuous flavour symmetry. The symmetry current typically lives in the anti-symmetric rank-2 tensor representation of the flavour symmetry. E. g., if the flavour symmetry is $\mathrm{U}(1)$, then the symmetry current has charge-0 ; if the flavour symmetry is $\mathrm{O}(3)$, then the symmetry current has spin-1 and is odd under the improper $\mathbb{Z}_{2}$ transformation ; if the flavour symmetry is $\mathrm{O}(n)$ ($n\geq 4$) or $\mathrm{SU}(n)$ ($n\geq 3$), then the symmetry current lives in the anti-symmetric rank-$2$ tensor representation.
@@
    where $\Box=\partial_{\mu}\partial^{\mu}$. For a spinning primary $\Phi^{\mu_{1}\dots\mu_{s}}$, its descendants have two forms :
    $\Box^{n}\partial^{\nu_{1}}\dots\partial^{\nu_{m}}\partial_{\rho_{1}}\dots\partial_{\rho_{k}}\Phi^{\rho_{1}\dots\rho_{k}\mu_{1}\dots\mu_{s-k}}\qquad(k=0,\dots,s,\quad n,m=0,1,\dots)$
@@
    $\Box^{n}\partial^{\nu_{1}}\dots\partial^{\nu_{m}}\partial_{\rho_{1}}\dots\partial_{\rho_{k}}\epsilon^{\sigma}{}_{\tilde{\mu}\tilde{\nu}}\partial^{\tilde{\nu}}\Phi^{\rho_{1}\dots\rho_{k}\tilde{\mu}\mu_{1}\dots\mu_{s-k-1}}\qquad(k=0,\dots,s-1,\quad n,m=0,1,\dots)$
@@
    For the second form, the fully anti-symmetric tensor $\epsilon$ alters the parity. For conserved currents only $m=0$ descendants exist due to the conservation $\partial_{\mu_{1}}\Phi^{\mu_{1}\dots\mu_{l}}=0$.
@@
Alternatively, one can define a cost function that depends on the tuning parameter and the speed of light and compares the scaling dimensions obtained from the fuzzy sphere and the prediction by conformal symmetry. E. g., for the Ising CFT, the tuning parameters are the pseudo-potentials $\{U_{i}\}$ and the transverse field $h$. The criteria for conformal symmetry we use include the stress tensor $T^{\mu\nu}$ and the descendants $\partial^{\mu}\sigma$, $\partial^{\mu}\partial^{\nu}\sigma$, $\Box\sigma$, $\partial^{\mu}\epsilon$. The cost function is the root-mean-square of the deviations of these criteria from the expectation of the conformal symmetry
$Q^{2}(\{U_{i}\},h,v;N_{m})=\frac{1}{N_{s}}\left[(\Delta_{T^{\mu\nu}}^{\text{(FS)}}-3)^{2}+(\Delta_{\partial^{\mu}\sigma}^{\text{(FS)}}-\Delta_{\sigma}^{\text{(FS)}}-1)^{2}\right.\\ \left.+(\Delta_{\partial^{\mu}\partial^{\nu}\sigma}^{\text{(FS)}}-\Delta_{\sigma}^{\text{(FS)}}-1)^{2}+(\Delta_{\Box\sigma}^{\text{(FS)}}-\Delta_{\sigma}^{\text{(FS)}}-1)^{2}+(\Delta_{\partial^{\mu}\epsilon}^{\text{(FS)}}-\Delta_{\epsilon}^{\text{(FS)}}-1)^{2}\right]$ (39)
@@
$\Delta_{\Phi}^{\text{(FS)}}(\{U_{i}\},h,v;N_{m})=\frac{E_{\Phi}-E_{0}}{v/R}.$ (40)
@@ 4.6 Local Observables
We have introduced how to determine the scaling dimensions from the energy spectrum. Beyond that, evaluating other CFT quantities requires realising local CFT operators on the fuzzy sphere. Any gapless local observables $\mathscr{O}(\mathbf{r})$ on the fuzzy sphere can be written as the linear combination of CFT operators that live in the same representation of flavour symmetry and parity [^9] The realisation of CFT operators in the microscopic model has been investigated in more detail in 2D in Ref. [145]. [2]
$\mathscr{O}(\mathbf{r},\tau)=\sum_{\alpha}\lambda_{\alpha}\Phi^{\text{(cyl.)}}_{\alpha}(\mathbf{r},\tau).$ (41)
Here special care should be taken for the CFT operator $\Phi^{\text{(cyl.)}}_{\alpha}(\mathbf{r},\tau)$ on the cylinder. A conformal transformation produces a scale factor $\Lambda(\mathbf{r})^{\Delta_{\Phi}}$ to a primary operator $\Phi$. The scale factor is $\Lambda(\mathbf{r})=r/R$ for the Weyl transformation from the flat space-time to the cylinder. Hence, [^10] We need to clarify some of the notations : $\mathscr{O}$ represents an operator in the microscopic model, and $\Phi$ represents a CFT operator ; the arguments $\Phi(\mathbf{r})$ or $\Phi(\mathbf{r},\tau)$ by default mean the operator is defined on a cylinder, and $\Phi(x)$ by default on flat space-time.
$\Phi^{\text{(cyl.)}}_{\alpha}(\mathbf{r},\tau)=\left(\frac{e^{\tau/R}}{R}\right)^{\Delta_{\Phi_{\alpha}}}\Phi_{\alpha}^{\text{(flat)}}(x).$ (42)
For descendants, certain other factors may be produced, but the conversion factors still scale with the radius of the sphere as $R^{-\Delta}$ where $\Delta$ is the scaling dimension of the descendants. For simplicity, hereafter we focus on the equal-time correlators with $\tau=0$, for which $\Phi^{\text{(cyl.)}}_{\alpha}(\mathbf{r})=R^{-\Delta_{\Phi_{\alpha}}}\Phi_{\alpha}^{\text{(flat)}}(x)$. The operator with larger system size decays faster when increasing system size.
The simplest local observable is the density operator defined in Eq. (14), and its decomposition into angular modes is given in Eqs. (16) and (17). From the CFT perspective, the density operators are the superpositions of scaling operators with corresponding quantum numbers, i. e. with the same representation under flavour symmetry and parity.
@@
$n^{x}(\mathbf{r})$ $=\lambda_{0}+\lambda_{\epsilon}\epsilon(\mathbf{r})+\lambda_{\partial^{\mu}\epsilon}\partial^{\mu}\epsilon(\mathbf{r})+\lambda_{T^{\mu\nu}}T^{\mu\nu}(\mathbf{r})+\dots$ $\epsilon_{\textrm{FS}}$ $=\frac{n^{x}-\lambda_{0}}{\lambda_{\epsilon}}+\dots$
$n^{z}(\mathbf{r})$ $=\lambda_{\sigma}\sigma(\mathbf{r})+\lambda_{\partial^{\mu}\sigma}\partial^{\mu}\epsilon(\mathbf{r})+\lambda_{\partial^{\mu}\partial^{\nu}\sigma}\partial^{\mu}\partial^{\nu}\sigma(\mathbf{r})+\dots$ $\sigma_{\textrm{FS}}$ $=\frac{n^{z}}{\lambda_{\sigma}}+\dots$ (43)
@@
We first consider the insertion of a single operator $\langle\Phi_{1}|\Phi_{2}(\mathbf{r})|\Phi_{3}\rangle$. It helps us produce the OPE coefficients. For the simplest example of three scalars,
$f_{\Phi_{1}\Phi_{2}\Phi_{3}}=\lim_{r_{\infty}\to\infty}r_{\infty}^{-2\Delta_{\Phi_{1}}}\langle\Phi_{1}(x_{\infty})\Phi_{2}(x)\Phi_{3}(0)\rangle_{\text{(flat)}}=\langle\Phi_{1}|\Phi_{2}^{\text{(flat)}}(x)|\Phi_{3}\rangle$ (44)
@@
After the Weyl transformation from the flat space-time to the cylinder, we obtain the expression on the fuzzy sphere
$f_{\Phi_{1}\Phi_{2}\Phi_{3}}=R^{\Delta_{\Phi_{2}}}\langle\Phi_{1}|\Phi^{(\text{(cyl.)})}_{2}(\mathbf{r})|\Phi_{3}\rangle.$ (47)
The UV realisation of $\Phi_{2}$ contains many other operators with different spins. By integrating the correlation function against different spherical harmonics, i. e. take the angular modes of the operator inserted
$\int\mathrm{d}^{2}\mathbf{r}\,\bar{Y}_{lm}(\hat{\mathbf{n}})\langle\Phi_{1}|\Phi_{2}(\mathbf{r})|\Phi_{3}\rangle=\langle\Phi_{1}|\Phi_{2,lm}|\Phi_{3}\rangle,$ (48)
we can filter out the sub-leading contributions with different spins. For the spinning operators, this also tells us about different OPE structures. By taking $\Phi_{3}=\mathbb{I}$, we can recover the 2-pt functions
@@
$\Phi_{2}(\mathbf{r})|0\rangle$ $=R^{-\Phi_{2}}\left[|\Phi_{2}\rangle+\lambda^{\prime}_{\mu}(\mathbf{r})|\partial^{\mu}\Phi_{2}\rangle+\lambda^{\prime\prime}(\mathbf{r})|\Box\Phi_{2}\rangle+\lambda^{\prime\prime}_{\mu\nu}(\mathbf{r})|\partial^{\mu}\partial^{\nu}\Phi_{2}\rangle\right].$ (49)
It is worth noting that acting a primary $\Phi_{2}(\mathbf{r})$ on the vacuum also produces various descendants in the multiplet.
@@
$\lambda_{0}=\frac{1}{\sqrt{4\pi}}\langle 0|n^{x}_{00}|0\rangle,\quad\lambda_{\epsilon}=\frac{R^{\Delta_{\epsilon}}}{\sqrt{4\pi}}\langle\epsilon|n^{x}_{00}|0\rangle,\quad\lambda_{\sigma}=\frac{R^{\Delta_{\sigma}}}{\sqrt{4\pi}}\langle\sigma|n^{z}_{00}|0\rangle.$ (50)
@@
$f_{\sigma\sigma\epsilon}$ $=R^{\Delta_{\sigma}}\langle\epsilon|\sigma(\mathbf{r})|\sigma\rangle=\frac{\langle\epsilon|n_{00}^{z}|\sigma\rangle}{\langle 0|n_{00}^{z}|\sigma\rangle}+\mathscr{O}(R^{-2})$
$=R^{\Delta_{\epsilon}}\langle\sigma|\epsilon(\mathbf{r})|\sigma\rangle=\frac{\langle\sigma|n^{x}_{00}|\sigma\rangle-\langle 0|n^{x}_{00}|0\rangle}{\langle\epsilon|n^{x}_{00}|0\rangle}+\mathscr{O}(R^{-(3-\Delta_{\epsilon})}).$ (51)
For the first line, the sub-leading contribution comes from the contribution of the descendant $\Box\sigma$ to $n_{00}^{z}$. As $\sigma(\mathbf{r})$ scales as $R^{-\Delta_{\sigma}}$ and $\Box\sigma(\mathbf{r})$ as $R^{-\Delta_{\sigma}-2}$,
@@
Here $c_{1}$ and $c^{\prime}_{1}$ are constant factors that represent the contribution of $\Box\sigma$ and do not scale with system size. Hence, the sub-leading contribution scales as $R^{-2}$. For the second line, the sub-leading contribution comes from the stress tensor $T^{\mu\nu}$. Similarly, the power of the scaling is the difference of the scaling dimension $R^{-(\Delta_{T^{\mu\nu}}-\Delta_{\epsilon})}=R^{-(3-\Delta_{\epsilon})}$.